# Explainable Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

### v1

@@ Abstract
In-situ sensing, in conjunction with learning models, presents a unique opportunity to address persistent defect issues in Additive Manufacturing (AM) processes. However, this integration introduces significant data privacy concerns, such as data leakage, sensor data compromise, and model inversion attacks, revealing critical details about part design, material composition, and machine parameters. Differential Privacy (DP) models, which inject noise into data under mathematical guarantees, offer a nuanced balance between data utility and privacy by obscuring traces of sensing data. However, the introduction of noise into learning models, often functioning as black boxes, complicates the prediction of how specific noise levels impact model accuracy. This study introduces the Differential Privacy-HyperDimensional computing (DP-HD) framework, leveraging the explainability of the vector symbolic paradigm to predict the noise impact on the accuracy of in-situ monitoring, safeguarding sensitive data while maintaining operational efficiency. Experimental results on real-world high-speed melt pool data of AM for detecting overhang anomalies demonstrate that DP-HD achieves superior operational efficiency, prediction accuracy, and robust privacy protection, outperforming state-of-the-art Machine Learning (ML) models. For example, when implementing the same level of privacy protection (with a privacy budget set at 1), our model achieved an accuracy of 94.43%, surpassing the performance of traditional models such as ResNet50 (52.30%), GoogLeNet (23.85%), AlexNet (55.78%), DenseNet201 (69.13%), and EfficientNet B2 (40.81%). Notably, DP-HD maintains high performance under substantial noise additions designed to enhance privacy, unlike current models that suffer significant accuracy declines under high privacy constraints.
Keywords Explainable AI  $\cdot$ Differential Privacy  $\cdot$ In-situ Sensing and Monitoring  $\cdot$ Smart Additive Manufacturing
@@ 1 Introduction
The integration of in-situ sensing and monitoring significantly advances agile defect detection, drastically reducing production time and costs by facilitating real-time adjustments to process parameters, thus enhancing the repeatability and reliability of AM processes [7]. Machine Learning (ML) models have demonstrated a significant enhancement in AM is extensively employed to take advantage of data to automatically diagnose printing status [8], failure modes [9], calibration [10], melting condition [11, 12], porosity detection [13], tensile property prediction [14], and surface roughness prediction [15], as well as perform feedback control [16, 17], model-based feed forward control [18, 19], and distortion compensation [20]. However, significant privacy risks arise due to the manufacturing sector’s vulnerability to cyberattacks. In 2022, the manufacturing sector experienced the highest rate of cyber incidents at 23.2%, surpassing even finance and insurance [21]. These cyberattacks, including ransomware and data breaches, have serious financial impacts, with average breach costs exceeding \$4.35 million. Small and medium-sized manufacturing firms are particularly vulnerable; in 2017, 61% of these firms reported cyberattacks, with median breach costs over \$60,000 [22]. Implementing in-situ sensing techniques exposes sensitive operational data, increasing the risk of unauthorized access. Model inversion attacks are particularly concerning, as adversaries exploit ML model vulnerabilities to extract confidential data about product specifications, process parameters, and machinery used in predictive analytics [23, 24].
@@
To mitigate privacy concerns in in-situ sensing, several privacy-preserving techniques are employed, including cryptographic methods, data anonymization, and Differential Privacy (DP). Cryptographic methods safeguard identities and manage data access, ensuring the confidentiality and integrity of sensitive information [25]. However, substantial computational power and higher bandwidth are often required, hindering the real-time processing essential for in-situ monitoring. Data anonymization involves removing identifying information from datasets prior to analysis [26], aiding in the management of complex, high-dimensional data, and reducing the risk of data breaches. Nevertheless, this technique results in the loss of valuable data granularity and does not fully guarantee privacy. The potential for re-identification remains high, as demonstrated by instances where anonymized Netflix data was re-identified by correlating it with external IMDb data [27].
@@
DP has emerged as an effective method for ensuring confidentiality in AM by strategically adding controlled noise to data without compromising data privacy [28, 29]. DP has been successfully applied in various industries, notably healthcare, to protect sensitive information against model inversion attacks [30, 31]. However, employing DP requires careful management of the trade-off between enhanced privacy and the performance of ML models. Higher levels of introduced noise can strengthen privacy protections but may also degrade algorithm precision and functionality. Thus, finding an optimal balance that ensures robust data protection while retaining process utility and efficiency is essential. The challenge is further compounded by the inherent "black box" nature of ML models, which lack transparency in how they derive decisions on noise level. This limited explainability poses significant challenges in achieving a balanced optimization of privacy and performance. Consequently, there is a critical need for advancements in explainable artificial intelligence to incorporate dynamics and enhance the transparency of ML models with DP in AM [32].
@@
To effectively predict the impact of DP noise on model accuracy, this paper introduces Differential Privacy-HyperDimensional computing (DP-HD). This explainable solution is designed for in-process monitoring, characterized by rapid query processing capabilities, making it ideal for scenarios requiring swift analysis while preserving both privacy and performance integrity. The effectiveness of DP-HD originates from its unique method of data handling, which leverages high-dimensional vector spaces for computation. This approach not only facilitates quick analysis but also allows for the precise definition of privacy levels and the prediction of model performance before any noise is added. Our contributions to this paper are as follows:
@@
    Propose explainable DP-HD to enhance the transparency of interactions between ML models and DP. The framework introduces a method for determining the optimal standard deviation for generating random basis vectors, ensuring an optimal trade-off between privacy and performance.
@@
    Evaluate the impact of noise on the framework’s memorization capability to demonstrate DP-HD’s ability to predict the effects of noise on performance and maintain high accuracy under various privacy settings.
@@
    Present an inference privacy mechanism that reduces the risk of decoding query hypervectors by eliminating low-variance dimensions, thus enhancing the security of the DP-HD model during the inference phase.
The structure of this paper is organized as follows: Section 2 discusses previous works related to in-situ sensing in AM and the preservation of privacy within these systems. Section 3 details the proposed methodology. Section 4 outlines the experimental setup and describes the design used in our case study, utilizing real-world AM data. Section 5 discusses the findings from the experiments. Finally, Section 6 concludes the paper by highlighting the limitations of current optimization methods for parameter selection in AM and suggesting potential avenues for future research.
@@ 2.1 In-situ Sensing and Monitoring
AM employs various sensing techniques for defect detection, including laser scanning [33], thermographic methods [34], radiographic [35], ultrasound [36], electromagnetic [37], and acoustic emission [38]. Visual imaging methods, in particular, are prized for their ability to automate the detection and evaluation of defects, making them critical for in-situ quality assurance in AM [39]. Commonly used equipment includes visible light cameras with specialized filters and fast shutters, or infrared cameras to monitor the AM process [40]. For example, Barua et al. utilized digital cameras with macro lenses to observe the laser metal deposition melt pool, creating a temperature map from RGB data through linear regression, which helps in identifying defects by highlighting variations in thermal conductivity [41].
@@
While some defect detection methods for AM processes support in-situ applications, many are unable to effectively identify small defects or lack reliable accuracy metrics [47]. ML models ensure product quality and enhance the efficiency of the printing process in AM [48], as they excel at extracting relevant features from complex datasets. ML has been widely applied in manufacturing contexts [49], highlighting its capability for defect detection in real-time, examining powder spreading, and optimizing process parameters [50]. Recent studies have successfully employed ML in various AM processes. For instance, defect detection in laser powder bed fusion has been improved using diverse ML and deep learning techniques, achieving accuracies above 80% [51, 52]. Moreover, ML and deep learning methods have been applied to quality monitoring in selective laser melting [53] and fused filament fabrication [54]. Convolutional Neural Networks (CNNs) have also been proposed to assess the quality of the powder bed fusion process by utilizing high-speed cameras for image data acquisition [55].
State-of-the-art ML models for image classification include ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. ResNet50 employs deep residual learning, allowing for the training of deeper networks through residual blocks with skip connections, which improve gradient flow and mitigate the vanishing gradient problem, achieving excellent results on complex tasks [56]. GoogLeNet, also known as Inception v1, uses inception modules to perform convolutions of various sizes alongside max pooling, optimizing computational resources and enhancing network depth and width without significant computational overhead [57]. AlexNet revitalized interest in CNNs with its deep convolutional layers, max-pooling, ReLU activations, dropout, and fully connected layers, drastically reducing error rates in image classification tasks [58]. DenseNet201 connects each layer to every other layer in a feed-forward manner, promoting feature reuse and reducing the number of parameters through dense connectivity, which supports effective feature propagation and efficiency in preserving feature information [59]. EfficientNet B2 builds on the base EfficientNet architecture by applying a compound scaling coefficient, balancing network depth, width, and resolution to enhance performance while being resource-efficient, making it ideal for high accuracy in constrained environments [60].
@@ 2.2 Privacy-Preserving in In-situ Sensing
To tackle privacy concerns, incorporating noise into data processing algorithms, as initially proposed by Dwork, is vital. This technique ensures that databases, even if they differ by just one record, remain indistinguishable when processed by the same algorithm, thus greatly enhancing privacy protection [29]. The intentional addition of noise effectively obscures data, preventing the leakage of sensitive information and guarding against model inversion attacks. By embedding DP into AM processes, manufacturers can protect confidential data while allowing for secure and privacy-aware applications of ML in monitoring and analysis [24, 61]. As AM technology progresses, incorporating DP and other privacy-preserving strategies will be essential in overcoming the sector’s cybersecurity issues. These methods not only ensure the privacy of individual data but also provide a safeguard against advanced cyber threats, maintaining the security and integrity of the manufacturing process. However, implementing these methods necessitates a delicate balance between increasing privacy and maintaining algorithmic performance. While higher noise levels enhance privacy, they can also diminish the precision and efficiency of the algorithm. Therefore, it is crucial to establish an optimal balance that provides robust data protection while preserving the process’s utility and effectiveness.
@@
The state-of-the-art method for applying DP to ML models is DP-SGD, a variation of the standard SGD algorithm. DP-SGD enhances privacy by introducing precisely calibrated noise to the gradients during the training process. This technique minimizes the impact of individual data entries by clipping gradients and adding noise, thereby masking their specific contributions to the final model parameters. DP-SGD provides measurable privacy assurances, essential in fields where data confidentiality is paramount [62].
Integrating DP into AM processes allows manufacturers to protect confidential data while enabling secure and privacy-conscious applications of ML for monitoring and analysis. However, introducing noise into ML models, especially those operating as black-box systems, complicates the understanding of the noise’s impact on model accuracy. This complexity makes it challenging to predict how specific noise levels will affect performance. To address this issue, we propose DP-HD as a novel solution to improve the explainability of interactions between ML models and DP. This approach aims to resolve the challenges of balancing privacy and performance, providing a clearer understanding of how DP modifications influence ML model outcomes.

### v3

# Explainable Differential Privacy-Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

@@ Abstract
Machine Learning (ML) models combined with in-situ sensing offer a powerful solution to address defect detection challenges in Additive Manufacturing (AM), yet this integration raises critical data privacy concerns, such as data leakage and sensor data compromise, potentially exposing sensitive information about part design and material composition. Differential Privacy (DP), which adds mathematically controlled noise to ML models, provides a way to balance data utility with privacy by concealing identifiable traces from sensor data. However, introducing noise into ML models, especially black-box Artificial Intelligence (AI) models, complicates the prediction of how noise impacts model accuracy. This study presents the Differential Privacy-Hyperdimensional Computing (DP-HD) framework, which leverages Explainable AI (XAI) and the vector symbolic paradigm to quantify noise effects on accuracy. By defining a Signal-to-Noise Ratio (SNR) metric, DP-HD assesses the contribution of training data relative to DP noise, allowing selection of an optimal balance between accuracy and privacy. Experimental results using high-speed melt pool data for anomaly detection in AM demonstrate that DP-HD achieves superior operational efficiency, prediction accuracy, and privacy protection. For instance, with a privacy budget set at 1, DP-HD achieves $94.43\%$ accuracy, outperforming state-of-the-art ML models. Furthermore, DP-HD maintains high accuracy under substantial noise additions to enhance privacy, unlike current models that experience significant accuracy declines under stringent privacy constraints.
Keywords Explainable Artificial Intelligence  $\cdot$ Hyperdimensional Computing  $\cdot$ Differential Privacy  $\cdot$ Additive Manufacturing Monitoring
@@ 1 Introduction
The integration of in-situ sensing and monitoring significantly advances agile defect detection, drastically reducing production time and costs by facilitating real-time adjustments to process parameters, thus enhancing the repeatability and reliability of AM processes [7]. Machine Learning (ML) models have significantly enhanced AM by leveraging data to automatically diagnose printing status [8], identify failure modes [9], calibrate systems [10], assess melting conditions [11, 12], detect porosity [13], predict tensile properties [14] and surface roughness [15], as well as to perform feedback control [16, 17], model-based feed-forward control [18, 19], and distortion compensation [20]. However, significant privacy risks arise due to the manufacturing sector’s vulnerability to cyberattacks. In 2022, the manufacturing sector experienced the highest rate of cyber incidents at 23.2%, surpassing even finance and insurance [21]. These cyberattacks, including ransomware and data breaches, have serious financial impacts, with average breach costs exceeding \$4.35 million. Small and medium-sized manufacturing firms are particularly vulnerable; in 2017, 61% of these firms reported cyberattacks, with median breach costs over \$60,000 [22]. Implementing in-situ sensing techniques exposes sensitive operational data, increasing the risk of unauthorized access. Model inversion attacks are particularly concerning, as adversaries exploit ML model vulnerabilities to extract confidential data about product specifications, process parameters, and machinery used in predictive analytics [23, 24].
@@
To mitigate privacy concerns in in-situ sensing, several privacy-preserving techniques are employed, including cryptographic methods, data anonymization, and Differential Privacy (DP). Cryptographic methods safeguard identities and manage data access, ensuring the confidentiality and integrity of sensitive information [25]. However, substantial computational power and higher bandwidth are often required, hindering the real-time processing essential for in-situ monitoring. Data anonymization involves removing identifying information from datasets prior to analysis [26], aiding in the management of complex, high-dimensional data, and reducing the risk of data breaches. Nevertheless, this technique results in the loss of valuable data granularity and does not fully guarantee privacy. The potential for re-identification remains high, as demonstrated by instances where anonymized Netflix data was re-identified by correlating it with external data [27].
@@
DP has emerged as an effective method for ensuring confidentiality in AM by strategically adding controlled noise to data to protect data privacy [28, 29]. DP has been successfully applied in various industries, notably healthcare, to protect sensitive information against model inversion attacks [30, 31]. However, employing DP requires careful management of the trade-off between enhanced privacy and the performance of ML models. Higher levels of noise can strengthen privacy protections but may also degrade algorithm precision and functionality. Thus, finding an optimal balance that ensures robust data protection while retaining process utility and efficiency is essential. A significant gap in existing approaches is compounded by the "black-box" nature of ML models, which obscures how DP noise impacts model accuracy, making it difficult to predict or fine-tune the privacy-accuracy balance. This lack of transparency and predictability in black-box models poses considerable limitations in optimizing DP applications effectively. Consequently, there is a critical need for advancements in Explainable Artificial Intelligence (XAI) that make the effects of DP noise on model accuracy more interpretable and quantifiable, particularly in privacy-sensitive AM applications [32].
@@
To effectively anticipate the impact of DP noise on model accuracy, this paper introduces the Differential Privacy-Hyperdimensional Computing (DP-HD) framework. This explainable approach is tailored for in-process monitoring, featuring rapid query response capabilities, making it ideal for applications requiring swift analysis while preserving both privacy and performance fidelity. The strength of DP-HD lies in its innovative data processing method, which utilizes high-dimensional vector spaces for computation. This design not only enables efficient analysis but also provides a way to define privacy levels precisely and predict model accuracy before adding noise, aided by the introduction of a Signal-to-Noise Ratio (SNR) metric to measure the balance between data contribution and DP noise. The primary contributions of this work are as follows:
@@
    We propose an explainable DP-HD framework to improve transparency in interactions between ML models and DP. The framework introduces a method for determining the optimal Standard Deviation (SD) for generating random basis vectors, aided by the use of an SNR metric to achieve an ideal balance between privacy protection and performance accuracy.
@@
    We assess the impact of noise on the framework’s memorization capacity, utilizing SNR to showcase DP-HD’s ability to predict noise effects on model accuracy across various privacy levels.
@@
    We present an inference privacy technique that mitigates the risk of decoding query hypervectors by removing low-variance dimensions, further enhancing the security of DP-HD during inference.
The structure of this paper is organized as follows: Section 2 discusses previous works related to in-situ sensing in AM and the preservation of privacy within these systems. Section 3 details the proposed methodology. Section 4 outlines the experimental setup and describes the design used in our case study, utilizing real-world AM data. Section 5 discusses the findings from the experiments. Section 6 provides an analysis of the implications of our results, examining the strengths, limitations, and broader impact of the proposed framework. Finally, Section 7 concludes the paper by highlighting the limitations of current optimization methods for parameter selection in AM and suggesting potential avenues for future research.
@@ 2.1 In-situ Sensing and Monitoring
AM employs various sensing techniques for defect detection, including laser scanning [33], thermographic methods [34], radiographic [35], ultrasound [36], electromagnetic [37], and acoustic emission [38]. Visual imaging methods, in particular, are prized for their ability to automate the detection and evaluation of defects, making them critical for in-situ quality assurance in AM [39]. Commonly used equipment includes visible light cameras with specialized filters and fast shutters, or infrared cameras to monitor the AM process [40]. For example, Barua et al. utilized digital cameras with macro lenses to observe the laser metal deposition melt pool, creating a temperature map from Red, Green, and Blue (RGB) data through linear regression, which helps in identifying defects by highlighting variations in thermal conductivity [41].
@@
While some defect detection methods for AM processes support in-situ applications, many are unable to effectively identify small defects or lack reliable accuracy metrics [47]. ML models ensure product quality and enhance the efficiency of the printing process in AM [48], as they excel at extracting relevant features from complex datasets. ML has been widely applied in manufacturing contexts [49], highlighting its capability for defect detection in real-time, examining powder spreading, and optimizing process parameters [50]. Recent studies have successfully employed ML in various AM processes. For instance, defect detection in Laser Powder Bed Fusion (LPBF) has been improved using diverse ML and deep learning techniques, achieving accuracies above 80% [51, 52]. Moreover, ML and deep learning methods have been applied to quality monitoring in selective laser melting [53] and fused filament fabrication [54]. Convolutional Neural Networks (CNNs) have also been proposed to assess the quality of the powder bed fusion process by utilizing high-speed cameras for image data acquisition [55].
State-of-the-art ML models for image classification include ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. ResNet50 employs deep residual learning, allowing for the training of deeper networks through residual blocks with skip connections, which improve gradient flow and mitigate the vanishing gradient problem, achieving excellent results on complex tasks [56]. GoogLeNet, also known as Inception v1, uses inception modules to perform convolutions of various sizes alongside max pooling, optimizing computational resources and enhancing network depth and width without significant computational overhead [57]. AlexNet revitalized interest in CNNs with its deep convolutional layers, max-pooling, Rectified Linear Unit (ReLU) activation functions, dropout, and fully connected layers, drastically reducing error rates in image classification tasks [58]. DenseNet201 connects each layer to every other layer in a feed-forward manner, promoting feature reuse and reducing the number of parameters through dense connectivity, which supports effective feature propagation and efficiency in preserving feature information [59]. EfficientNet B2 builds on the base EfficientNet architecture by applying a compound scaling coefficient, balancing network depth, width, and resolution to enhance performance while being resource-efficient, making it ideal for high accuracy in constrained environments [60].
@@ 2.2 Privacy-Preserving in In-situ Sensing
Incorporating noise into algorithms, as originally proposed by Dwork, is essential for privacy. This technique ensures that databases, even if they differ by just one record, remain indistinguishable when processed by the same algorithm, thus greatly enhancing privacy protection [29]. The intentional addition of noise effectively obscures data, preventing the leakage of sensitive information and guarding against model inversion attacks. By embedding DP into AM processes, manufacturers can protect confidential data while allowing for secure and privacy-aware applications of ML in monitoring and analysis [24, 61]. As AM technology progresses, incorporating DP and other privacy-preserving strategies will be essential in overcoming the sector’s cybersecurity issues. These methods not only ensure the privacy of individual data but also provide a safeguard against advanced cyber threats, maintaining the security and integrity of the manufacturing process. However, implementing these methods necessitates a delicate balance between increasing privacy and maintaining algorithmic performance. While higher noise levels enhance privacy, they can also diminish the precision and efficiency of the algorithm. Therefore, it is crucial to establish an optimal balance that provides robust data protection while preserving the process’s utility and effectiveness.
@@
The state-of-the-art method for applying DP to ML models is Differentially Private Stochastic Gradient Descent (DP-SGD), a variation of the standard Stochastic Gradient Descent (SGD) algorithm. DP-SGD enhances privacy by introducing precisely calibrated noise to the gradients during the training process. This technique minimizes the impact of individual data entries by clipping gradients and adding noise, thereby masking their specific contributions to the final model parameters. DP-SGD provides measurable privacy assurances, essential in fields where data confidentiality is paramount [62].
2.3 Explainable AI
Integrating DP into AM processes enables manufacturers to secure sensitive data, supporting privacy-centric ML applications for monitoring and analysis. However, introducing noise into ML models, especially those operating as black-box systems, complicates the interpretation of how noise impacts model accuracy. Traditional XAI approaches, reviewed extensively in [63], primarily focus on improving interpretability in various fields but often lack the mechanisms to handle DP-specific challenges, such as the predictable impact of DP noise on model accuracy. This limitation highlights a gap in existing XAI methods, particularly under stringent privacy constraints, thereby underscoring the need for more specialized frameworks like DP-HD.
To address this gap, the DP-HD framework we propose employs an innovative approach centered around an SNR metric. This metric quantifies the contribution of training data relative to DP noise, establishing a structured foundation for balancing privacy and accuracy. Inspired by advanced XAI techniques found in recent applications like crop recommendation systems [64], DP-HD integrates an SNR-based interpretation layer to enhance transparency. This approach allows the framework to both monitor the effect of DP noise on model predictions and reliably estimate accuracy across privacy levels. By offering this quantifiable insight, DP-HD extends beyond conventional XAI limitations, providing a robust mathematical solution that clarifies the impact of DP noise in high-dimensional AM models while ensuring data privacy.
Table 1: Summary of XAI Techniques: Advantages and Limitations
XAI Technique Advantages Limitations
Attribution Methods [65] Provide clear relevance of input features to model decisions; easy to interpret as post-hoc explanations Depend on model internals, making them less suitable for black-box models
Saliency Methods (e.g., InputXGradient [66], GuidedBackpropagation [67], IntegratedGradients [68]) Efficiently highlight relevant features in neural networks using gradient-based approaches Require access to network internals; limited in handling black-box scenarios
Perturbation-based Methods (e.g., Occlusion [69]) Model-agnostic; can be applied to any architecture Resource-intensive as they involve systematic input modifications; limited interpretability in complex models
Surrogate Models (e.g., LIME [70], SHAP [71]) Provide local explanations; SHAP offers mathematical grounding through game theory Computationally expensive, especially with high-dimensional data; may provide overly simplified explanations in complex models
Prototype [72], Patch [73], and Concept-based Methods [74] Enable interpretable visual or semantic concepts linked to model predictions Limited by interpretability of visual or semantic features; do not address privacy concerns effectively
Within the broader field of XAI, a variety of techniques have been developed to make ML models more interpretable. Attribution methods, widely used for their adaptability and clarity, assess the contribution of input features to model predictions, particularly through post-hoc analysis methods [65]. Saliency methods [75] were among the first to apply backpropagation for explaining model predictions, with extensions like InputXGradient [66] and others (GuidedBackpropagation [67] and IntegratedGradients [68]) further refining this approach. These gradient-based techniques provide insight by accessing model internals, a feature unavailable in black-box contexts.
In contrast, perturbation-based methods like Occlusion [69] are model-agnostic, meaning they do not require access to specific model architectures. Such methods work by systematically altering inputs to observe the effect on model predictions. Other approaches, like Local Interpretable Model-Agnostic Explanations (LIME) [70], generate local surrogate models to explain predictions, while Shapley Additive Explanations (SHAP) [71] apply game theory principles to offer mathematically grounded insights into feature importance.
Beyond attribution-based methods, other XAI techniques, including prototype-based [72], patch-based [73], and concept-based methods [74], offer additional ways to interpret complex models. However, while these traditional XAI methods provide valuable interpretability, they often fall short of addressing DP noise effects on model accuracy, a critical aspect of balancing privacy and utility in high-stakes applications like AM.
The DP-HD framework offers a structured approach to balancing privacy and accuracy by introducing an SNR-based metric, making it well-suited for sensitive applications that require strict data privacy protections. While developed within the context of AM, the principles behind DP-HD can also be extended to other privacy-sensitive environments, such as healthcare, financial services, and government data management, where secure handling of confidential information is paramount. In these domains, where sensitive data like patient records, financial transactions, and personal identifiers must be protected, DP-HD’s ability to quantify and predict noise impact on model accuracy provides an advanced solution. This generalizability underscores the DP-HD framework’s potential as a versatile tool for privacy-preserving ML across various sectors, ensuring robust protection without sacrificing model performance. Table 1 summarizes the discussed XAI methods, highlighting their respective advantages and limitations.